# Quality and Quantity of Machine Translation References for Automated Metrics

### 02 Jan 2024 v1

@@ 5 Conclusion
We showed that the quality of references is important for accurate automated machine translation metrics. The relationships is not straightforward: translatologists’ translations, despite being the peak human translation quality, are not the best references. Rather, it is the standard commercial professional translation that work best for current metrics. The improvements apply both for string-matching metrics as well as to parametric model-based ones. Taking the average over multiple references provides the biggest benefit, with diminishing returns after 7 references. We also provided a heuristic-based algorithm for finding a good configuration of references given a budget, which surpasses optimizing solely for quantity or quantity.
@@
The dataset size prevents system-level investigations. Because there is little point in evaluating segments that are easy to translate, a follow-up approach could prioritize difficult-to translate segments. This is used by Isabelle et al. (2017) for creating a challenge set. Future works should extend this research by asking the question how many segments are needed to fulfill a certain desideratum, such as effect size or metric accuracy.
Limitations
We note the limitation of using only a small dataset and a single language direction due to the costs of creating multiple rounds of high-quality references. However, we are convinced the results hold in other scenarios because the effect directions are the same across multiple metrics and setups.
Input: Source segments , levels , cost function , utility function ,
tradeoff hyperparameter , temperature , budget .
Output: Assignment .
Note: Figure 3 contains a patience mechanism, spans and uses .
Sample a segment to add.
Sample where to move from.
Appendix A Budget Allocation Algorithm
We formalize the problem with a segment cost for a reference on level and the utility . The levels might correspond to translation vendors which have costs and qualities. In Section 3.4, we assumed and . Given a set of source sentences , the goal is to divvy up the segments into different levels. The selection should maximize performance of a particular metric on a number of systems but needs to be under a fixed budget , i.e. . In our setup, to preserve fair comparison, each segment needs to have at least one reference. This is because the smaller the testset, the easier it is to achieve higher but spurious correlations. Therefore, . The formalization explicitly allows for parts of the testset to be translated multiple times.
We present a heuristic algorithm with pseudocode in Algorithm 1 and explanatory illustration of the two operations in Figure 4. The algorithm continually applies one of the two operations until they can either no longer be applied or the budget is reached. The algorithm will always terminate because Add increases the cost and utility and Move increases the utility by at least and , respectively.
In addition to the metric correlations in Figure 3, we show the average number of references per source segment in Figure 5. With focus on quality, each segment has fewer references.
Appendix B Qualitative Analysis
In Table 6 we show a single source segment, system translation and multiple references and the metric scores. BLEU ranges from 0 to 100 and both maximums are almost achieved just with a different reference. The best human translation led to the lowest BLEU score because of a translation shift. This is not surprising because BLEU operates on the surface-level. Unexpectedly, a similar thing happens also with Comet, which uses a distributed semantic representation of the segments. This shows that parametric model-based metrics are not robust to changes in references.
@@ Appendix B Qualitative Analysis
Appendix C Implementation Details
BLEU, ChrF and TER used SacreBleu (Post 2018, v:2.3.1) at the segment-level. For the trainable metrics, we use publicly available models.
BLEU: #:1|c:mixed|e:yes|tok:13a|s:exp
ChrF: #:1|c:mixed|e:yes|nc:6|nw:0|s:no
TER: #:1|c:lc|t:tercom|nr:no|pn:yes|as:no
COMET: Unbabel/wmt20-comet-da
COMET: Unbabel/wmt22-comet-da
BLEURT: BLEURT20
| Metric | R1-R1 | R2-R2 | R3-R3 | R4-R4 | |
Layman PE |
BLEU | 0.019 | 0.007 | 0.009 | 0.010 |
| ChrF | 0.027 | 0.015 | 0.016 | 0.019 | |
| TER | 0.026 | 0.015 | 0.013 | 0.014 | |
| COMET | 0.016 | 0.011 | 0.010 | 0.009 | |
| COMET | 0.008 | 0.006 | 0.007 | 0.005 | |
| BLEURT | 0.017 | 0.015 | 0.011 | 0.010 | |
Student PE |
BLEU | 0.010 | 0.001 | 0.004 | 0.001 |
| ChrF | 0.011 | 0.001 | 0.000 | 0.004 | |
| TER | 0.003 | 0.001 | 0.005 | 0.002 | |
| COMET | 0.010 | 0.003 | 0.002 | 0.002 | |
| COMET | 0.002 | 0.001 | 0.000 | 0.002 | |
| BLEURT | 0.021 | 0.022 | 0.005 | 0.004 | |
Professional PE |
BLEU | 0.035 | 0.011 | 0.007 | 0.000 |
| ChrF | 0.040 | 0.010 | 0.008 | 0.004 | |
| TER | 0.023 | 0.014 | 0.003 | 0.002 | |
| COMET | 0.016 | 0.006 | 0.000 | 0.005 | |
| COMET | 0.008 | 0.002 | 0.003 | 0.005 | |
| BLEURT | 0.027 | 0.022 | 0.005 | 0.001 |
@@ Acknowledgements
We extend our gratitude to Yasmin Moslem, Isabella Lai, Theia Vogel, and Blaka Sokołowska who proofread this paper in their free time. Ondřej Bojar received funding from Ministry of Education, Youth and Sports of the Czech Republic LM2018101 LINDAT/CLARIAH-CZ and the 19-26934X grant of the Czech Science Foundation (NEUREM3).

### 07 Mar 2024 v4

@@ 6. Conclusion
We showed that the quality of references is important for accurate automatic machine translation metrics. The relationship is not straightforward: translatologists’ translations, despite being the peak translation quality, are not the best references. Rather, it is the standard commercial professional translations that work best for current metrics. The trend applies to both string-matching metrics as well as to parametric model-based ones. Taking the average over multiple references provides the biggest benefit, with diminishing returns after 7 references. We also provided a heuristic-based algorithm for finding a good configuration of references given a budget, which surpasses optimizing solely for quantity or quantity.
@@
The dataset size prevents system-level investigations. Because there is little point in evaluating segments that are easy to translate, a follow-up approach could prioritize difficult-to translate segments. This is used by Isabelle et al. (2017) for creating a challenge set. Future works should quantify the references quality and ask how many segments are needed to fulfill a certain desideratum, such as effect size or metric accuracy.
Limitations.
We note the limitation of using a small dataset and a single language translation direction due to the costs of creating multiple rounds of high-quality references. We are convinced the results hold in other scenarios as the effect directions are the same across multiple metrics and setups.
![[Uncaptioned image]](x20.png)
Figure 5: Average number of references per one segment allocated by Algorithm 1 with (top) and (bottom).
@@
| BLEU | COMET | Reference |
| 4 | -42 | Sony a Disney točí třetí film o Spidermanovi |
| 4 | -33 | Sony a Disney točí třetí film o Spider-Manovi |
| 8 | -9 | Sony a Disney pracují na třetím filmu o Spider-Manovi |
| 8 | -4 | Sony a Disney pokračují v práci na třetím filmu o Spider-Manovi |
| 8 | 1 | Sony a Disney opět pracují na třetím filmu o Spider-Manovi |
| 8 | 15 | Sony a Disney spolupracují na třetím filmu o Spider-Manovi |
| 4 | 16 | Sony a Disney budou spolupracovat při natáčení třetího filmo o Spider-manovi |
| 8 | 28 | Sony a Disney opět spolupracují na třetím filmu o Spider-Manovi |
| 8 | 30 | Sony a Disney budou spolupracovat na třetím filmu o Spider-Manovi |
| 8 | 35 | Sony a Disney budou opět spolupracovat na třetím filmu o Spider-Manovi |
| 17 | 52 | Disney bude znovu spolupracovat se společností Sony na třetím filmu Spider-Man |
| 10 | 64 | Disney bude se Sony dál pracovat na třetím filmu se Spider-Manem |
| 50 | 73 | Disney bude spolupracovat se Sony na třetím sólovém filmu o Spider-Manovi |
| 75 | 99 | Disney se vrací, bude spolupracovat se Sony na třetím filmu o Spider-Manovi |
| 78 | 106 | Disney se vrací, bude spolupracovat se Sony na třetím sólovém filu Spider-Man. |
| 79 | 108 | Disney se vrací, bude spolupracovat se Sony na třetím Spider-Man filmu |
| 100 | 121 | Disney se vrací, bude spolupracovat se Sony na třetím sólovém Spider-Man filmu |
@@ Acknowledgements
We extend our gratitude to Yasmin Moslem, Isabella Lai, Theia Vogel, Blanka Sokołowska, Raj Dabre, Tom Kocmi, and Farhan Samir, who proofread this paper in their free time. Ondřej Bojar received funding from Ministry of Education, Youth and Sports of the Czech Republic LM2018101 LINDAT/CLARIAH-CZ and the 19-26934X grant of the Czech Science Foundation (NEUREM3).
Bibliographical References
[…]
- Bojar et al. (2010) Ondřej Bojar, Kamil Kos, and David Mareček. 2010. Tackling sparse data issue in machine translation evaluation. In Proceedings of the ACL 2010 Conference Short Papers, 86–91, Uppsala, Sweden. Association for Computational Linguistics.
- Bojar et al. (2013) Ondřej Bojar, Christian Buck, Chris Callison-Burch, Christian Federmann, Barry Haddow, Philipp Koehn, Christof Monz, Matt Post, Radu Soricut, and Lucia Specia. 2013. Findings of the 2013 Workshop on Statistical Machine Translation. In Proceedings of the Eighth Workshop on Statistical Machine Translation, 1–44, Sofia, Bulgaria. Association for Computational Linguistics.
- Castilho et al. (2018) Sheila Castilho, Stephen Doherty, Federico Gaspari, and Joss Moorkens. 2018. Approaches to human and machine translation quality assessment. Translation quality assessment: From principles to practice.
- Daems and Macken (2020) Joke Daems and Lieve Macken. 2020. Post-editing human translations and revising machine translations: Impact on efficiency and quality. In Translation Revision and Post-editing, 50–70. Routledge.
- Finch et al. (2004) Andrew M Finch, Yasuhiro Akiba, and Eiichiro Sumita. 2004. How does automatic machine translation evaluation correlate with human scoring as the number of reference translations increases? In LREC.
- Fomicheva (2017) Marina Fomicheva. 2017. The role of human reference translation in machine translation evaluation. Ph.D. thesis, Universitat Pompeu Fabra.
- Fomicheva et al. (2020) Marina Fomicheva, Lucia Specia, and Francisco Guzmán. 2020. Multi-hypothesis machine translation evaluation. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics.
- Freitag et al. (2021a) Markus Freitag, George Foster, David Grangier, Viresh Ratnakar, Qijun Tan, and Wolfgang Macherey. 2021a. Experts, Errors, and Context: A Large-Scale Study of Human Evaluation for Machine Translation. Transactions of the Association for Computational Linguistics, 9:1460–1474.
- Freitag et al. (2020) Markus Freitag, David Grangier, and Isaac Caswell. 2020. BLEU might be guilty but references are not innocent. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 61–71. Association for Computational Linguistics.
- Freitag et al. (2023) Markus Freitag, Nitika Mathur, Chi-kiu Lo, Eleftherios Avramidis, Ricardo Rei, Brian Thompson, Tom Kocmi, Frederic Blain, Daniel Deutsch, Craig Stewart, Chrysoula Zerva, Sheila Castilho, Alon Lavie, and George Foster. 2023. Results of WMT23 metrics shared task: Metrics might be guilty but references are not innocent. In Proceedings of the Eighth Conference on Machine Translation, 578–628. Association for Computational Linguistics.
- Freitag et al. (2022) Markus Freitag, Ricardo Rei, Nitika Mathur, Chi-kiu Lo, Craig Stewart, Eleftherios Avramidis, Tom Kocmi, George Foster, Alon Lavie, and André F. T. Martins. 2022. Results of WMT22 metrics shared task: Stop using BLEU – neural metrics are better and more robust. In Proceedings of the Seventh Conference on Machine Translation, 46–68. Association for Computational Linguistics.
- Freitag et al. (2021b) Markus Freitag, Ricardo Rei, Nitika Mathur, Chi-kiu Lo, Craig Stewart, George Foster, Alon Lavie, and Ondřej Bojar. 2021b. Results of the WMT21 metrics shared task: Evaluating metrics with expert-based human evaluations on TED and news domain. In Proceedings of the Sixth Conference on Machine Translation, 733–774. Association for Computational Linguistics.
- González-Rubio et al. (2012) Jesús González-Rubio, Daniel Ortiz-Martínez, and Francisco Casacuberta. 2012. Active learning for interactive machine translation. In Proceedings of the 13th Conference of the European Chapter of the Association for Computational Linguistics, 245–254. Association for Computational Linguistics.
- Graham et al. (2016) Yvette Graham, Timothy Baldwin, Alistair Moffat, and Justin Zobel. 2016. Can machine translation systems be evaluated by the crowd alone. Natural Language Engineering, FirstView:1–28.
- Haffari et al. (2009) Gholamreza Haffari, Maxim Roy, and Anoop Sarkar. 2009. Active learning for statistical phrase-based machine translation. In Proceedings of Human Language Technologies: The 2009 Annual Conference of the North American Chapter of the Association for Computational Linguistics, 415–423. Association for Computational Linguistics.
- Hamon and Mostefa (2008) Olivier Hamon and Djamel Mostefa. 2008. The impact of reference quality on automatic MT evaluation. In Coling 2008: Companion volume: Posters, 39–42.
- Isabelle et al. (2017) Pierre Isabelle, Colin Cherry, and George Foster. 2017. A challenge set approach to evaluating machine translation. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 2486–2496. Association for Computational Linguistics.
- Khayrallah et al. (2020) Huda Khayrallah, Brian Thompson, Matt Post, and Philipp Koehn. 2020. Simulated multiple reference training improves low-resource machine translation. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 82–89. Association for Computational Linguistics.
- Kloudová et al. (2023) Věra Kloudová, David Mraček, Ondřej Bojar, and Martin Popel. 2023. Možnosti a meze tvorby tzv. optimálních referenčních překladů: po stopách „překladatelštiny“ v profesionálních překladech zpravodajských textů. Slovo a slovesnost, 84(2):122–156.
- Kocmi et al. (2023) Tom Kocmi, Eleftherios Avramidis, Rachel Bawden, Ondřej Bojar, Anton Dvorkovich, Christian Federmann, Mark Fishel, Markus Freitag, Thamme Gowda, Roman Grundkiewicz, Barry Haddow, Philipp Koehn, Benjamin Marie, Christof Monz, Makoto Morishita, Kenton Murray, Makoto Nagata, Toshiaki Nakazawa, Martin Popel, Maja Popović, and Mariya Shmatova. 2023. Findings of the 2023 conference on machine translation (WMT23): LLMs are here but not quite there yet. In Proceedings of the Eighth Conference on Machine Translation, 1–42, Singapore. Association for Computational Linguistics.
- Kocmi et al. (2021) Tom Kocmi, Christian Federmann, Roman Grundkiewicz, Marcin Junczys-Dowmunt, Hitokazu Matsushita, and Arul Menezes. 2021. To ship or not to ship: An extensive evaluation of automatic metrics for machine translation. In Proceedings of the Sixth Conference on Machine Translation, 478–494. Association for Computational Linguistics.
- Kocmi et al. (2024) Tom Kocmi, Vilém Zouhar, Christian Federmann, and Matt Post. 2024. Navigating the metrics maze: Reconciling score magnitudes and accuracies.
- Krubiński et al. (2021) Mateusz Krubiński, Erfan Ghadery, Marie-Francine Moens, and Pavel Pecina. 2021. Just ask! evaluating machine translation by asking and answering questions. In Proceedings of the Sixth Conference on Machine Translation, 495–506, Online. Association for Computational Linguistics.
- Lommel (2016) Arle Lommel. 2016. Blues for BLEU: Reconsidering the validity of reference-based MT evaluation. Translation Evaluation: From Fragmented Tools and Data Sets to an Integrated Ecosystem, 63.
- Macháček and Bojar (2014) Matouš Macháček and Ondřej Bojar. 2014. Results of the WMT14 metrics shared task. In Proceedings of the Ninth Workshop on Statistical Machine Translation, 293–301, Baltimore, MD, USA. Association for Computational Linguistics.
- Madnani et al. (2008) Nitin Madnani, Philip Resnik, Bonnie J. Dorr, and Richard Schwartz. 2008. Are multiple reference translations necessary? investigating the value of paraphrased reference translations in parameter optimization. In Proceedings of the 8th Conference of the Association for Machine Translation in the Americas, 143–152.
- Mathur et al. (2020) Nitika Mathur, Johnny Tian-Zheng Wei, Markus Freitag, Qingsong Ma, and Ondřej Bojar. 2020. Results of the WMT20 metrics shared task. In Fifth Conference on Machine Translation - Proceedings of the Conference, 688–725, Stroudsburg, PA, USA. Association for Computational Linguistics, Association for Computational Linguistics.
- Mendonça et al. (2023) Vânia Mendonça, Ricardo Rei, Luísa Coheur, and Alberto Sardinha. 2023. Onception: Active Learning with Expert Advice for Real World Machine Translation. Computational Linguistics, 49(2):325–372.
- Mi et al. (2020) Chenggang Mi, Lei Xie, and Yanning Zhang. 2020. Improving adversarial neural machine translation for morphologically rich language. IEEE Transactions on Emerging Topics in Computational Intelligence, (4):417–426.
- Ott et al. (2018) Myle Ott, Michael Auli, David Grangier, and Marc’Aurelio Ranzato. 2018. Analyzing uncertainty in neural machine translation. In Proceedings of the 35th International Conference on Machine Learning, volume 80 of Proceedings of Machine Learning Research, 3956–3965. PMLR.
- Papineni et al. (2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002. Bleu: a method for automatic evaluation of machine translation. In Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics, 311–318. Association for Computational Linguistics.
- Popović (2015) Maja Popović. 2015. chrF: character n-gram F-score for automatic MT evaluation. In Proceedings of the Tenth Workshop on Statistical Machine Translation, 392–395. Association for Computational Linguistics.
- Qin and Specia (2015) Ying Qin and Lucia Specia. 2015. Truly exploring multiple references for machine translation evaluation. In Proceedings of the 18th Annual Conference of the European Association for Machine Translation. European Association for Machine Translation.
- Rei et al. (2022) Ricardo Rei, José G. C. de Souza, Duarte Alves, Chrysoula Zerva, Ana C Farinha, Taisiya Glushkova, Alon Lavie, Luisa Coheur, and André F. T. Martins. 2022. COMET-22: Unbabel-IST 2022 submission for the metrics shared task. In Proceedings of the Seventh Conference on Machine Translation, 578–585. Association for Computational Linguistics.
- Rei et al. (2021) Ricardo Rei, Ana C Farinha, Chrysoula Zerva, Daan van Stigt, Craig Stewart, Pedro Ramos, Taisiya Glushkova, André F. T. Martins, and Alon Lavie. 2021. Are references really needed? unbabel-IST 2021 submission for the metrics shared task. In Proceedings of the Sixth Conference on Machine Translation, 1030–1040. Association for Computational Linguistics.
- Rei et al. (2020) Ricardo Rei, Craig Stewart, Ana C Farinha, and Alon Lavie. 2020. COMET: A neural framework for MT evaluation. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 2685–2702. Association for Computational Linguistics.
- Sellam et al. (2020) Thibault Sellam, Dipanjan Das, and Ankur Parikh. 2020. BLEURT: Learning robust metrics for text generation. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, 7881–7892. Association for Computational Linguistics.
- Shi and Huang (2020) Zhouxing Shi and Minlie Huang. 2020. Robustness to modification with shared words in paraphrase identification. In Findings of the Association for Computational Linguistics: EMNLP 2020, 164–171. Association for Computational Linguistics.
- Snover et al. (2006) Matthew Snover, Bonnie Dorr, Rich Schwartz, Linnea Micciulla, and John Makhoul. 2006. A study of translation edit rate with targeted human annotation. In Proceedings of the 7th Conference of the Association for Machine Translation in the Americas: Technical Papers, 223–231. Association for Machine Translation in the Americas.
- Specia and Shah (2018) Lucia Specia and Kashif Shah. 2018. Machine translation quality estimation: Applications and future perspectives. Translation quality assessment: from principles to practice, 201–235.
- van der Wees et al. (2017) Marlies van der Wees, Arianna Bisazza, and Christof Monz. 2017. Dynamic data selection for neural machine translation. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 1400–1410. Association for Computational Linguistics.
- Vernikos et al. (2022) Giorgos Vernikos, Brian Thompson, Prashant Mathur, and Marcello Federico. 2022. Embarrassingly easy document-level MT metrics: How to convert any pretrained metric into a document-level metric. In Proceedings of the Seventh Conference on Machine Translation (WMT), 118–128. Association for Computational Linguistics.
- Zhang and Vogel (2004) Ying Zhang and Stephan Vogel. 2004. Measuring confidence intervals for the machine translation evaluation metrics. In Proceedings of the 10th Conference on Theoretical and Methodological Issues in Machine Translation of Natural Languages.
- Zhang and Vogel (2010) Ying Zhang and Stephan Vogel. 2010. Significance tests of automatic machine translation evaluation metrics. Machine Translation, 24:51–65.
- Zheng et al. (2018) Renjie Zheng, Mingbo Ma, and Liang Huang. 2018. Multi-reference training with pseudo-references for neural translation and text generation. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 3188–3197. Association for Computational Linguistics.
- Zouhar et al. (2021) Vilém Zouhar, Martin Popel, Ondřej Bojar, and Aleš Tamchyna. 2021. Neural machine translation quality and post-editing performance. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, 10204–10214. Association for Computational Linguistics.
- Zouhar et al. (2023) Vilém Zouhar, Věra Kloudová, Martin Popel, and Ondřej Bojar. 2023. Evaluating optimal reference translations.